OpenStudy (anonymous):

A square has the same area as an equilateral triangle. What is the ratio of the perimeter of the equilateral triangle to the perimeter of the square?

OpenStudy (anonymous):

Area of Equilateral Triangle is how much can you tell??

OpenStudy (anonymous):

Do you know I am asking you??

OpenStudy (anonymous):

A Square with side a has Area = \(s^2\) Area of Equilateral Triangle = \(\frac{\sqrt{3}}{4} \times a^2\)

OpenStudy (anonymous):

Then I think you know the answer too..

OpenStudy (anonymous):

I got 3^.25a : 2s, but the answer is 3^.75a : 2s

OpenStudy (anonymous):

\[\large s^2 = \frac{\sqrt{3}}{4} \times a^2 \implies s = \frac{(3)^{\frac{1}{4}}a}{4}\]

OpenStudy (anonymous):

You forgot the square root of 4.

OpenStudy (anonymous):

Now take the ratio: \[\frac{3a}{4s} \implies \frac{3 \times 2}{4 \cdot (3)^{\frac{1}{4}}}\] Solve this..

OpenStudy (anonymous):

\[\sqrt[4]{3} = (3)^{\frac{1}{4}}\]

OpenStudy (anonymous):

3^.75/2

OpenStudy (anonymous):

That is the ratio..

OpenStudy (anonymous):

Why are you taking s and a there?? They will be cancel out on substitution..

OpenStudy (anonymous):

*get cancelled..

OpenStudy (anonymous):

I'm still confused.

OpenStudy (anonymous):

Yeah that is 2 instead of 4.. My mistake there..

OpenStudy (anonymous):

Where?? You can freely tell me..

OpenStudy (anonymous):

Got ??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Good..